You have a purified alcohol dehydrogenase from liver. You take 50 micro liters of this solution, and add 950 micro liters of buffer. The absorbance of the diluted solution is 0.542.

A) what is the ADH concentration in the undiluted sample (in mg/ml)? The extinction coefficient for a 0.1% solution of ADH at 280 Bm is 1.83.

B) what is the ADH concentration in micro M? (The molecular weight of ADH is 140,000 Da)
Answer #1

* To know the ADH concentration, the absorbance of the given sample is divided by the extinction coefficient. 0.542/1.83 = 0.296/0.1% solution. = 2.96/ for 1% solution.The original sample is 1ml, so there is 2.96mg/ml of protein.

* IN terms of micromoles, the protein concentration is 29617 micro moles